# ToMA: Token Merge with Attention for Diffusion Models

### v2

@@ Abstract
Diffusion models excel in high-fidelity image generation but face scalability limits due to transformers’ quadratic attention complexity. Plug-and-play token reduction methods like ToMeSD and ToFu reduce FLOPs by merging redundant tokens in generated images but rely on GPU-inefficient operations (e.g., sorting, scattered writes), introducing overheads that negate theoretical speedups when paired with optimized attention implementations (e.g., FlashAttention). To bridge this gap, we propose Token Merge with Attention (ToMA), an off-the-shelf method that redesigns token reduction for GPU-aligned efficiency, with three key contributions: 1) a reformulation of token merge as a submodular optimization problem to select diverse tokens; 2) merge/unmerge as an attention-like linear transformation via GPU-friendly matrix operations; and 3) exploiting latent locality and sequential redundancy (pattern reuse) to minimize overhead. ToMA reduces SDXL/Flux generation latency by 24%/23% (with DINO $\Delta<$ 0.07), outperforming prior methods. This work bridges the gap between theoretical and practical efficiency for transformers in diffusion.
@@ 1 Introduction
Diffusion models (Ho et al., 2020; Song et al., 2021; Dhariwal & Nichol, 2021) have revolutionized high-fidelity image generation. Yet, their reliance on transformer architectures—notably in U-ViT (Bao et al., 2023) and DiT (Peebles & Xie, 2023)—introduces a bottleneck: the quadratic complexity of self-attention scales prohibitively with token counts, exacerbating latency across denoising steps.
Efforts to accelerate transformers broadly fall into two categories: attention optimization (e.g., Flash Attention (Dao, 2023), xformers (Lefaudeux et al., 2022)) and token reduction. While there is extensive research on token reduction for discriminative tasks, including Token Merge (Bolya et al., 2023), AdaViT (Yin et al., 2022b), and later improvements like Diffrate (Chen et al., 2023), these approaches often employ irreversible token pruning or merging, which is unsuitable for generative tasks.
Generative architectures like diffusion models impose stricter constraints: token counts must be restored after merging (i.e., “unmerging”) to preserve spatial consistency for iterative refinement. Earlier attempts, including ToMeSD (Bolya & Hoffman, 2023), Token Pruning (Kim et al., 2022), and ToFu (Kim et al., 2023) adapt token reduction to this setting but suffer from a critical flaw: their merge/prune (and unmerge/fill) operations rely on GPU-inefficient primitives (e.g., sorting, scattered memory writes). This problem intensifies when paired with highly optimized attention implementations like those in the diffusers framework (von Platen et al., 2022).
@@ 2 Related Work
Vision Transformers (Dosovitskiy et al., 2021) face computational challenges due to quadratic attention complexity. Recent efforts to mitigate this fall into four categories: compact architectures (e.g., Swin Transformer (Liu et al., 2021a), PVT (Wang et al., 2021)), pruning strategies like X-Pruner (Yu & Xiang, 2023), knowledge distillation such as DeiT (Touvron et al., 2021), and post-training quantization (Liu et al., 2021b). Complementary efforts also explore combined techniques (Papa et al., 2024) to unify these paradigms. While effective, most solutions require retraining and remain inherently token-centric. In contrast, ToMA introduces training-free token merging, operating orthogonally, thus enabling seamless integration without conflict.
@@
Most learned token reduction involves training auxiliary models to assess the importance of tokens in the input data. DynamicViT (Rao et al., 2021) employs a lightweight MLP to generate pruning masks based on input token features. A-ViT (Yin et al., 2022a) computes halting probabilities using specific channels of token features to determine the necessity of further processing. Whereas these methods often require additional training for the auxiliary modules, our approach is directly applicable, offering a more generalizable solution.
@@
Heuristic token reduction strategies can be applied to ViTs without additional training. Our method also falls into this category, making the approaches below natural baselines for comparison. Adaptive Token Sampling (ATS) (Fayyaz et al., 2022) keeps tokens most similar to the class token, which limits its use in pixel-level generation tasks where a class token is absent. Token Downsampling (ToDo) (Smith et al., 2024) downsamples only the key–value in attention, skipping queries, which limits acceleration and causes fine-grained detail loss due to uniform spatial pooling. Token Merging for Stable Diffusion (ToMeSD) (Bolya & Hoffman, 2023) forms source–destination pairs within fixed or randomly tiled regions and greedily matches them based on similarity, performing unweighted merging followed by a simple unmerge that copies the destination embedding back to each source position. ToFu (Kim et al., 2023) builds on ToMeSD by dynamically deciding, for each layer, whether to merge or prune tokens according to a linearity test, thereby combining the benefits of both operations.
@@ 3.2 Submodularity
A submodular function (Fujishige, 2005) is a set function $f:2^{{\mathbb{V}}}\rightarrow\mathbb{R}$ defined over subsets of the ground set $\mathbb{V}$}. It satisfies the diminishing returns property, which states that the marginal gain of adding a new element $v$ to the set decreases as the context set grows. Mathematically:
@@
An intuitive approach is the greedy algorithm (Alg. 1), which guarantees a (11/e)(1-1/e)-approximation of the optimal solution (Nemhauser et al., 1978). Starting with $\mathbb{A}$$\emptyset$, the algorithm iteratively selects the element with the highest marginal gain until the constraint $|{\mathbb{A}}|=k$ is reached.
@@ 4.1 Submodular-Based Destination Selection
While there exist more advanced submodular maximization methods such as the Lazier-than-Lazy Greedy algorithm (Mirzasoleiman et al., 2015), their reliance on operations like random subset sampling introduces irregular memory access patterns that are inefficient on GPUs. Therefore, the standard greedy approach strikes a practical balance between solution quality and hardware efficiency.
@@ 5.1 Descriptions
We evaluate ToMA on two of the most widely used diffusion models: the UNet-based SDXL-base and the DiT-based Flux.1-dev, to generate $1024\times 1024$ images, using the Diffusers framework. Importantly, ToMA is architecture-agnostic and can be readily extended to other diffusion models (e.g., SD2, SD3.5). Prompts are drawn from the GEMRec dataset (Guo et al., 2024) and ImageNet–1K names of the classes (Deng et al., 2009).
@@
To assess image quality, we use CLIP-T, DINO, and FID (Radford et al., 2021; Caron et al., 2021; Heusel et al., 2017). CLIP-T measures semantic alignment between images and prompts via cosine similarity between image-text embeddings. DINO measures perceptual consistency by comparing visual features between the original generated image and its counterpart produced with the merge method applied. FID (Fréchet Inception Distance) quantifies distributional similarity to real images based on Inception-V3 feature statistics. For FID CLIP-T and DINO in the main experiments shown in this section, we generate 3,000 images and compute scores against ground-truths from ImageNet–1K. For the ablation experiments listed in Appendix F, we generate images from 50 prompts with three random seeds each and report the average. FID is omitted on GEMRec due to the lack of paired images. Inference latency is reported as the median wall-clock time over 100 runs.
@@ 5.2 Results
Table 3 presents a comparative evaluation of ToMA (our default method), alongside other token reduction strategies on the SDXL model. Among all methods, ToMA achieves the fastest generation time—up to 28.5% improvement—while maintaining competitive perceptual quality, making it the most balanced choice. Although ToMe (Token Merge) delivers slightly better image quality, as reflected in CLIP-T and DINO scores, it suffers from severe latency due to its complex token selection mechanism, which relies on GPU-inefficient operations (e.g., sort). ToFu (Token Fusion), while faster than ToMe with lower overhead thanks to its blended strategy that incorporates token pruning, exhibits unstable visual quality—some generations appear acceptable, but others are heavily degraded, especially at higher reduction ratios. The high FID values confirm this observation. Lastly, ToDo (Token Downsampling) offers moderate gains in both speed and fidelity but demonstrates noticeable distributional shifts, as indicated by its elevated FID score. Hwang et al. discuss the possible cause that the noise reduction due to downsampling creates a difference in the signal-to-noise ratio that leads to suboptimal performance if applied directly (2024). Overall, these results highlight ToMA as the most robust and well-rounded token merging strategy for efficient high-resolution image generation.

### v3

@@ Abstract
Diffusion models excel in high-fidelity image generation but face scalability limits due to transformers’ quadratic attention complexity. Plug-and-play token reduction methods like ToMeSD and ToFu reduce FLOPs by merging redundant tokens in generated images but rely on GPU-inefficient operations (e.g., sorting, scattered writes), introducing overheads that negate theoretical speedups when paired with optimized attention implementations (e.g., FlashAttention). To bridge this gap, we propose Token Merge with Attention (ToMA), an off-the-shelf method that redesigns token reduction for GPU-aligned efficiency, with three key contributions: 1) a reformulation of token merge as a submodular optimization problem to select diverse tokens; 2) merge/unmerge as an attention-like linear transformation via GPU-friendly matrix operations; and 3) exploiting latent locality and sequential redundancy (pattern reuse) to minimize overhead. ToMA reduces SDXL/Flux generation latency by 24%/23% (with DINO $\Delta<$ 0.07), outperforming prior methods. This work bridges the gap between theoretical and practical efficiency for transformers in diffusion. Code available at github.com/WenboLuu/ToMA.
@@ 1 Introduction
Diffusion models (ho2020denoising; song2021scorebased; dhariwal2021diffusion) have revolutionized high-fidelity image generation. Yet, their reliance on transformer architectures—notably in U-ViT (bao2023all) and DiT (peebles2023scalable)—introduces a bottleneck: the quadratic complexity of self-attention scales prohibitively with token counts, exacerbating latency across denoising steps.
Efforts to accelerate transformers broadly fall into two categories: attention optimization (e.g., Flash Attention (dao2023flashattention), xformers (xFormers2022)) and token reduction. While there is extensive research on token reduction for discriminative tasks, including Token Merge (Bolya2023-ne), AdaViT (yin2022vit), and later improvements like Diffrate (chen2023diffrate), these approaches often employ irreversible token pruning or merging, which is unsuitable for generative tasks.
Generative architectures like diffusion models impose stricter constraints: token counts must be restored after merging (i.e., “unmerging”) to preserve spatial consistency for iterative refinement. Earlier attempts, including ToMeSD (Bolya-tomesd), Token Pruning (tokenprune), and ToFu (ToFU) adapt token reduction to this setting but suffer from a critical flaw: their merge/prune (and unmerge/fill) operations rely on GPU-inefficient primitives (e.g., sorting, scattered memory writes). This problem intensifies when paired with highly optimized attention implementations like those in the diffusers framework (von-platen-etal-2022-diffusers).
@@ 2 Related Work
Vision Transformers (dosovitskiy2021imageworth16x16words) face computational challenges due to quadratic attention complexity. Recent efforts to mitigate this fall into four categories: compact architectures (e.g., Swin Transformer (liu2021swin), PVT (wang2021pyramid)), pruning strategies like X-Pruner (yu2023x), knowledge distillation such as DeiT (touvron2021training), and post-training quantization (liu2021posttrainingquantizationvisiontransformer). Complementary efforts also explore combined techniques (Papa_2024) to unify these paradigms. While effective, most solutions require retraining and remain inherently token-centric. In contrast, ToMA introduces training-free token merging, operating orthogonally, thus enabling seamless integration without conflict.
@@
Most learned token reduction involves training auxiliary models to assess the importance of tokens in the input data. DynamicViT (Rao2021-fm) employs a lightweight MLP to generate pruning masks based on input token features. A-ViT (Yin2022-hr) computes halting probabilities using specific channels of token features to determine the necessity of further processing. Whereas these methods often require additional training for the auxiliary modules, our approach is directly applicable, offering a more generalizable solution.
@@
Heuristic token reduction strategies can be applied to ViTs without additional training. Our method also falls into this category, making the approaches below natural baselines for comparison. Adaptive Token Sampling (ATS) (Fayyaz2022-fz) keeps tokens most similar to the class token, which limits its use in pixel-level generation tasks where a class token is absent. Token Downsampling (ToDo) (smith2024todo) downsamples only the key–value in attention, skipping queries, which limits acceleration and causes fine-grained detail loss due to uniform spatial pooling. Token Merging for Stable Diffusion (ToMeSD) (Bolya-tomesd) forms source–destination pairs within fixed or randomly tiled regions and greedily matches them based on similarity, performing unweighted merging followed by a simple unmerge that copies the destination embedding back to each source position. ToFu (ToFU) builds on ToMeSD by dynamically deciding, for each layer, whether to merge or prune tokens according to a linearity test, thereby combining the benefits of both operations.
@@ 3.2 Submodularity
A submodular function (fujishige2005submodular) is a set function $f:2^{{\mathbb{V}}}\rightarrow\mathbb{R}$ defined over subsets of the ground set $\mathbb{V}$}. It satisfies the diminishing returns property, which states that the marginal gain of adding a new element $v$ to the set decreases as the context set grows. Mathematically:
@@
An intuitive approach is the greedy algorithm (Alg. 1), which guarantees a (11/e)(1-1/e)-approximation of the optimal solution (nemhauser1978analysis). Starting with $\mathbb{A}$$\emptyset$, the algorithm iteratively selects the element with the highest marginal gain until the constraint $|{\mathbb{A}}|=k$ is reached.
@@ 4.1 Submodular-Based Destination Selection
While there exist more advanced submodular maximization methods such as the Lazier-than-Lazy Greedy algorithm (mirzasoleiman2015lazier), their reliance on operations like random subset sampling introduces irregular memory access patterns that are inefficient on GPUs. Therefore, the standard greedy approach strikes a practical balance between solution quality and hardware efficiency.
@@ 5.1 Descriptions
We evaluate ToMA on two of the most widely used diffusion models: the UNet-based SDXL-base and the DiT-based Flux.1-dev, to generate $1024\times 1024$ images, using the Diffusers framework. Importantly, ToMA is architecture-agnostic and can be readily extended to other diffusion models (e.g., SD2, SD3.5). Prompts are drawn from the GEMRec dataset (GEMRec) and ImageNet–1K names of the classes (imagenet1k).
@@
To assess image quality, we use CLIP-T, DINO, and FID (CLIP; dino; fid). CLIP-T measures semantic alignment between images and prompts via cosine similarity between image-text embeddings. DINO measures perceptual consistency by comparing visual features between the original generated image and its counterpart produced with the merge method applied. FID (Fréchet Inception Distance) quantifies distributional similarity to real images based on Inception-V3 feature statistics. For FID CLIP-T and DINO in the main experiments shown in this section, we generate 3,000 images and compute scores against ground-truths from ImageNet–1K. For the ablation experiments listed in Appendix F, we generate images from 50 prompts with three random seeds each and report the average. FID is omitted on GEMRec due to the lack of paired images. Inference latency is reported as the median wall-clock time over 100 runs.
@@ 5.2 Results
Table 3 presents a comparative evaluation of ToMA (our default method), alongside other token reduction strategies on the SDXL model. Among all methods, ToMA achieves the fastest generation time—up to 28.5% improvement—while maintaining competitive perceptual quality, making it the most balanced choice. Although ToMe (Token Merge) delivers slightly better image quality, as reflected in CLIP-T and DINO scores, it suffers from severe latency due to its complex token selection mechanism, which relies on GPU-inefficient operations (e.g., sort). ToFu (Token Fusion), while faster than ToMe with lower overhead thanks to its blended strategy that incorporates token pruning, exhibits unstable visual quality—some generations appear acceptable, but others are heavily degraded, especially at higher reduction ratios. The high FID values confirm this observation. Lastly, ToDo (Token Downsampling) offers moderate gains in both speed and fidelity but demonstrates noticeable distributional shifts, as indicated by its elevated FID score. Hwang et al. discuss the possible cause that the noise reduction due to downsampling creates a difference in the signal-to-noise ratio that leads to suboptimal performance if applied directly (hwang2024upsample). Overall, these results highlight ToMA as the most robust and well-rounded token merging strategy for efficient high-resolution image generation.